# Jones Wenzl projectors in Verma modules

### v1

@@ Abstract
We construct special idempotents in $\mathrm{End}_{U_{q}(\mathfrak{sl}_{2})}(M(\mu_{1})\otimes\cdots\otimes M(\mu_{n}))$ like the Jones Wenzl projector where $M(\mu_{i})$ is Verma module whose highest weight is $\mu_{i}$.
@@ 1 Introduction
Temperley Lieb algebra is introduced in [TemperleyLieb] to solve the problems of statistical physics. Furthermore, Wenzl introduced the special idempotents in Temperley Lieb algebra called Jones Wenzl idempotent in [Wenzl]. The Jones Wenzl idempotent plays an important role in topology. Murakami and Murakami constructed knot invariants using the Jones Wenzl idempotent in [MurakamiMurakami]. The knot invariants extend the Jones polynomial obtained in [Jones]. It is known that Temperley Lieb algebra is isomorphic to the endomorphism algebra of tensor products of $2$-dimensional irreducible representation over $U_{q}(\mathfrak{sl}_{2})$ in [AndersenLehrerZhang]. Then it is important to consider the idempotents in endomorphism algebras of tensor products of $U_{q}(\mathfrak{sl}_{2})$ representations. In [LacabanneTubbenhauerVaz], the structure of the endomorphism algebra of tensor products of Verma modules over $U_{q}(\mathfrak{sl}_{2})$ is determined using Howe duality-like method. However, the special idempotent of the endomorphism algebra like the Jones Wenzl idempotent has not been constructed. In this article, we construct special idempotents in $\mathrm{End}_{U_{q}(\mathfrak{sl}_{2})}(M(\mu_{1})\otimes\cdots\otimes M(\mu_{n}))$ like Jones Wenzl projector. More precisely, we obtain the following.
@@ 2 Representation of $U_{q}(\mathfrak{sl}_{2})$
In this section, we define the quantum algebra $U_{q}(\mathfrak{sl}_{2})$ and the representations.
@@
Verma module $M(\mu)$ over $\mathbb{C}(q,q^{\mu})$ has a basis $\{v_{0},v_{1},\cdots\}$ called an induced basis and satisfies the following relations for the generators of $U_{q}(\mathfrak{sl}_{2})$.
@@ 3 Endomorphism algebras of $V_{1}^{\otimes n}$ and $M(\mu_{1})\otimes\cdots\otimes M(\mu_{n})$
$EF_{\mu,\lambda}v_{k}$ $=\sum_{j=1}^{k}q^{-(k-j)(\mu-j)}\begin{bmatrix}k\\ j\end{bmatrix}\frac{\prod_{i=0}^{j-1}[\mu-i]\prod_{i=0}^{k-j-1}[\lambda-i]}{\prod_{i=0}^{k-1}[\mu+\lambda-i]}q^{\lambda-2(k-j)}[j]v_{j-1,k-j}$

### v2

@@ Abstract
We construct special idempotents in $\mathrm{End}_{U_{q}(\mathfrak{sl}_{2})}(M(\mu_{1})\otimes\cdots\otimes M(\mu_{n}))$ like the Jones Wenzl projector where $M(\mu_{i})$ is Verma module whose highest weight is $\mu_{i}$ and is complex number except non-negative integer.
@@ 1 Introduction
Temperley Lieb algebra is introduced in [TemperleyLieb] to solve the problems of statistical physics. Furthermore, Wenzl introduced the special idempotents in Temperley Lieb algebra called Jones Wenzl idempotent in [Wenzl]. The Jones Wenzl idempotent plays an important role in topology. Murakami and Murakami constructed knot invariants using the Jones Wenzl idempotent in [MurakamiMurakami]. The knot invariants extend the Jones polynomial obtained in [Jones]. It is known that Temperley Lieb algebra is isomorphic to the endomorphism algebra of tensor products of $2$-dimensional irreducible representation over $U_{q}(\mathfrak{sl}_{2})$ in [AndersenLehrerZhang]. Then it is important to consider the idempotents in endomorphism algebras of tensor products of $U_{q}(\mathfrak{sl}_{2})$ representations. In [LacabanneTubbenhauerVaz], the structure of the endomorphism algebra of tensor products of Verma modules over $U_{q}(\mathfrak{sl}_{2})$ is determined using Howe duality-like method. However, the special idempotent of the endomorphism algebra like the Jones Wenzl idempotent has not been constructed. In this article, we construct special idempotents in $\mathrm{End}_{U_{q}(\mathfrak{sl}_{2})}(M(\mu_{1})\otimes\cdots\otimes M(\mu_{n}))$ like Jones Wenzl projector where $\mu_{i}$ is complex number without non-negative integer. More precisely, we obtain the following.
@@ 2 Representation of $U_{q}(\mathfrak{sl}_{2})$
In this section, we define the quantum algebra $U_{q}(\mathfrak{sl}_{2})$ and the representations. Hereinafter, suppose that $q$ is generic.
@@
Let $\mu$ be a complex number without non-negative integer. Verma module $M(\mu)$ over $\mathbb{C}(q,q^{\mu})$ has a basis $\{v_{0},v_{1},\cdots\}$ called an induced basis and satisfies the following relations for the generators of $U_{q}(\mathfrak{sl}_{2})$.
@@ 3 Endomorphism algebras of $V_{1}^{\otimes n}$ and $M(\mu_{1})\otimes\cdots\otimes M(\mu_{n})$
$EF_{\mu,\lambda}v_{k}=$
$\sum_{j=1}^{k}q^{-(k-j)(\mu-j)}\begin{bmatrix}k\\ j\end{bmatrix}\frac{\prod_{i=0}^{j-1}[\mu-i]\prod_{i=0}^{k-j-1}[\lambda-i]}{\prod_{i=0}^{k-1}[\mu+\lambda-i]}q^{\lambda-2(k-j)}[j]v_{j-1,k-j}$